# Relativistic Jets and Winds in Radio-Identified Supermassive Black Hole Binary Candidates

### v1

@@ 1 Introduction
The discovery of a nHz gravitational wave (GW) background by the global network of pulsar timing arrays (PTAs) provides strong evidence for the ubiquity of supermassive black hole binaries (SMBHBs) throughout the universe (Burke-Spolaor et al., 2019; Agazie et al., 2023a, b; EPTA Collaboration et al., 2023; Reardon et al., 2023; Xu et al., 2023). Direct evidence for such binaries includes 4C+37.11 (Maness et al., 2004; Romani et al., 2014; Bansal et al., 2017; Surti et al., 2024), whose black holes have projected separation $7.3$ pc, as well as the quasi-periodic blazar OJ 287 (Sillanpaa et al., 1988; Komossa et al., 2022). However, despite the promise of searching for periodicity in active galactic nuclei (AGN) light curves as a discovery method, statistically strong SMBHB candidates have proven elusive (Vaughan et al., 2016; Liu et al., 2019).
@@
PKS 2131, a BL Lac object at $z=1.283$, has shown coherent $P=1739.8\pm 17.4$ day periodicity over 47.9 years in radio (Ren et al., 2021a; O’Neill et al., 2022; Kiehlmann et al., 2025). PKS J0805, a flat-spectrum radio quasar (FSRQ) at $z=1.388$ (Healey et al., 2007, 2008), has shown a $P=1240.7\pm 4.6$ day periodicity over 14 years of radio observation (Ren et al., 2021b; de la Parra et al., 2025). Recent observations have revealed the same periodicity across the spectrum, from cm to optical wavelengths, and possibly even gamma-rays (Kiehlmann et al., 2025; Hincks et al., 2025).
Motivated by the observations of PKS 2131 and PKS J0805, we present a model for jetted SMBHB systems. We invoke a sub-relativistic wind to collimate the jet launched from one of the two black holes in the binary. The wind, dragged around by the orbital motion, creates a conical-helical channel through which the relativistic flow travels, producing sinusoidal aberration. [^1] Note that, at radii much greater than the orbital radius, it is the orbital velocity that determines the wind direction. The jet direction consequently varies in quadrature with the displacement and so the jet kinematics is similar to that induced by jet precession. This work enhances the simple model presented in O’Neill et al. (2022) and Kiehlmann et al. (2025), in which the relativistic jet propagates ballistically, requiring coherent variation over $\sim 100$ pc scales. We detail the model in Sec. 2. We interpret the observations of PKS 2131 and PKS J0805 and apply our model in Sec. 3. In Sec. 4, we consider the prospects of future emission line and GW observations, before concluding in Sec. 5.
@@ 2 Flux variation from Disk-wind Collimated Jets
Figure 1: A visualization of our model. (Top) An SMBHB composed of the primary black hole H and a secondary black hole S. H orbits with speed $v_{H}$, carries around a disk (D) and launches a jet (J) along the orbital angular momentum direction (L). The disk produces a wind (W) with speed $v_{w}<c$, that collimates the jet. (Middle) A geometric illustration of the conical jet along the $\hat{z}$ axis with opening angle $\theta_{j}$ confined by the wind. The jet wind interface has velocity $\vec{\beta}_{w}=\vec{v}_{w}/c$. We show a unit vector for $\hat{r}$ associated with angular position ($\chi$, $\psi$) as well as the observer viewing direction $\hat{n}$. The orbital motion with velocity $\vec{\beta}_{H}=\vec{v}_{H}/c$ will distort the jet into a conical helix. (Bottom) A 2D slice through the wind-collimated helical jet. The dashed line in the center denotes the sinuous path of relativistic flow moving through the center of the channel. The minimum radii $r_{\min}$ which contribute at two frequencies are labeled.
In our model, the SMBHB has period $P$ and is composed of primary mass $M_{1}$ and secondary mass $M_{2}$. We envision one black hole (either the primary or secondary) launching a jet. If the other black hole also launches a jet, we can add a second jet whose orbital motion is half a period out of phase. The main observational feature will be additional higher harmonics. We imagine a sub-relativistic collimating wind, which confines a synchrotron emitting relativistic jet with velocity $\beta_{j}$ and associated bulk Lorentz factor $\Gamma_{j}$. We suppose the interface between the jet and wind has effective speed $\beta_{w}\lesssim 0.9$ (i.e. lying between the sub-relativistic wind and highly relativistic jet speeds). In this model, the wind plays a different role from the sheath in Sobacchi et al. (2017) because it does not contribute to the radiation. The wind serves primarily to confine the jet, while the radiation comes from inside the jet.
The sub-relativistic wind may be a hydromagnetic flow produced from an extended disk (e.g. Blandford & Payne, 1982; Emmering et al., 1992; Globus & Levinson, 2016) or a jet sheath (e.g. Pelletier & Roland, 1989; Sol et al., 1989; Sobacchi et al., 2017). We assume the inertia of the wind-jet complex is predominantly carried by the wind. Consequently, we imagine that the relativistic jet direction is completely determined by the wind-jet interface as in the middle panel of Fig. 1. The orbital motion will add to the effective wind velocity, distorting the direction sinusoidally, forming a conical helix as in the bottom panel of Fig. 1. Functionally, this model is similar to those presented in Sobacchi et al. (2017) and O’Neill et al. (2022), except that the helix shape and propagation speed are set by the wind-jet interface rather than a ballistically propagating jet.
@@
The inclusion of the wind in our model adds a lever on the emission region locations at different frequencies. If the jet were to propagate ballistically with $\Gamma_{j}$, phase offsets in the observed sinusoidal light curves at different observing frequencies seen in PKS 2131 and PKS J0805 would suggest emission regions separated by $\sim 100$ pc (Kiehlmann et al., 2025). These large separations would require the helical structure to remain coherent over $\gtrsim 100$ helical wavelengths $\lambda_{h}=\beta_{w}cP$ (where $\beta_{w}=\beta_{j}$ if the relativistic flow is ballistic). With the jet channel propagating at a non-ultra-relativistic $\beta_{w}$, the emission regions can be separated by $\lesssim 30\lambda_{h}$.
@@ 2.1 Basic Jet Model
We suppose that the jet is confined within a cone of constant opening angle $\theta_{j}$. The observer direction makes an angle $i$ with the cone axis. The jet velocity is radial and constant throughout the jet, so that $\vec{\beta}_{j}=\beta_{j}\hat{r}$, where $\hat{r}$ is a radial unit vector associated with a particular angular position within the open jet (see middle panel of Fig. 1). We assume that the isotropic relativistic electron distribution function in the comoving jet (primed) frame is $n^{\prime}(\gamma)=K^{\prime}\gamma^{-p}$ for $\gamma_{e,\rm min}<\gamma<\gamma_{e,\rm max}$. The associated electron pressure is $P^{\prime}_{e}=\frac{1}{3}K^{\prime}m_{e}c^{2}\int_{\gamma_{e,\mathrm{m}in}}^{\gamma_{e,\mathrm{m}ax}}\gamma^{-p}d\gamma$. We normalize the pressure using the power in relativistic electrons $L_{e}=4\pi P^{\prime}_{e}(\Gamma_{j}\theta_{j}r)^{2}c$. Likewise, the comoving magnetic field strength $B^{\prime}$, assumed to be mostly transverse and disordered, is normalized by the electromagnetic power $L_{m}=(B^{\prime}\Gamma_{j}\theta_{j}r)^{2}c/4$.
@@
The synchrotron emissivity in the observer frame at frequency $\nu$ will be
@@
$\begin{split}j_{\nu,\Omega}=&D^{\frac{3+p}{2}}j^{\prime}_{\nu,\Omega}\\ =&D^{\frac{3+p}{2}}\frac{\sqrt{3\pi}}{16\pi^{2}m_{e}c^{2}}e^{3}\frac{K^{\prime}}{p+1}\left(\frac{2\pi m_{e}c}{3e}\right)^{-\frac{p-1}{2}}B^{\prime\frac{p+1}{2}}\\ &\times{\mathrm{\Gamma}}\left(\frac{p}{4}+\frac{19}{12}\right){{\Gamma}}\left(\frac{p}{4}-\frac{1}{12}\right)\\ &\times{{\Gamma}}\left(\frac{p}{4}+\frac{5}{4}\right){\Gamma}\left(\frac{p}{4}+\frac{7}{4}\right)^{-1}\nu^{-\alpha},\end{split}$ (2)
@@
where $j^{\prime}_{\nu,\Omega}$ is the emissivity in the relativistic flow frame, $\Gamma$ denotes the Gamma function (not to be confused with jet bulk Lorentz factor $\Gamma_{j}$), and $\alpha=(p-1)/2$ is the spectral index (Rybicki & Lightman, 1979). The absorption coefficient in the observer frame is
@@
$\begin{split}\mu=&D^{\frac{p+2}{2}}\mu^{\prime}\\ =&D^{\frac{p+2}{2}}\frac{\sqrt{3\pi}}{32\pi^{2}m_{e}^{2}c^{2}}\left(\frac{3e}{2\pi m_{e}c}\right)^{\frac{p}{2}}K^{\prime}B^{\prime\frac{p+2}{2}}\\ &\times\Gamma\left(\frac{p}{4}+\frac{22}{12}\right)\Gamma\left(\frac{p}{4}+\frac{2}{12}\right)\\ &\times\Gamma\left(\frac{p}{4}+\frac{6}{4}\right)\Gamma\left(\frac{p}{4}+\frac{8}{4}\right)^{-1}\nu^{-\frac{p+4}{2}},\end{split}$ (3)
@@
where $\mu^{\prime}$ denotes the absorption coefficient in the flow frame. For radio emitting particles in $B\sim 1$ mG fields, the synchrotron cooling times are much longer than the orbital period, so adiabatic losses dominate.
@@
Orbital motion causes major variation in $D$ and consequently flux density variation due to coherent aberration across the jet. To allow coherent helical motion, hydromagnetic waves must be able to cross the jet on a timescale comparable to the curvature radius of the wind-collimated jet (i.e. $\Gamma_{j}\theta_{j}\lesssim 1/2\pi\beta_{w}$). In practice, the jet kinematics is likely significantly more complicated and variable. We discuss some of these complications in Appendix A. The underlying “clock” should not exhibit frequency noise.
@@ 2.2 Kinematics
Due to the orbital motion and collimation by the wind, a radial vector $\hat{r}$ in a particular direction within the open jet will be redirected
@@
where $\hat{r}_{0}$ is the unperturbed radial normal vector, which corresponds to a fixed direction in the black hole frame. Note that $\vec{\beta}_{H}$ will introduce a sinusoidal perturbation to $\hat{r}$. At a particular radius $r$ from the black hole, $\vec{\beta}_{H}$ in eq. 4 is delayed in phase from the instantaneous orbital velocity by $r/\lambda_{h}$.
@@ 2.2.1 Amplitude of Flux Density Variations
in the observer frame. We express the line of sight direction as
@@
where $i$ the angle between the jet axis at its base and $\hat{n}$. The middle panel of Fig. 1 shows these vectors labeled on the jet.
@@ 2.2.2 Linear Polarization
Given that the magnetic field is assumed to be transverse [^2] Observational justification for our assumption of a toroidal magnetic field is given in Kiehlmann et al. (2025), as the polarization in PKS 2131 is observed to be along the large-scale jet direction., the electric vector should be along $\hat{r}$. The electric vector polarization angle (EVPA) and polarization degree (PD) will vary because the angle between $\hat{r}$ and $\hat{n}$ changes. The polarization properties discussed here will be for one specific position ($r,\chi,\psi$) in the jet. Different $(r,\chi,\psi)$ positions will have the same general polarization pattern but be slightly out of phase, leading to some decoherence.
@@
$\vec{e}_{obs}=\frac{\hat{r}-\left(\hat{r}\cdot\hat{n}\right)\hat{n}}{\sqrt{1-\left(\hat{r}\cdot\hat{n}\right)^{2}}},$ (18)
while the observed PD is
$\Pi_{\rm obs}=\Pi_{\max}\left[1-\left(\hat{r}\cdot\hat{n}\right)^{2}\right],$ (19)
where $\Pi_{\max}$ is maximum PD allowed given the level of turbulence. We decompose eq. 18 and eq. 19 into a constant and varying component

### v2

@@ Abstract
Radio jets (1347) – Active galactic nuclei (16) – Blazars (164) – Black holes (162)
@@ 1 Introduction
The discovery of a nHz gravitational wave (GW) background by the global network of pulsar timing arrays (PTAs) provides strong evidence for the ubiquity of supermassive black hole binaries (SMBHBs) throughout the universe (2019A&ARv..27....5B; 2023ApJ...951L...8A; 2023ApJ...952L..37A; 2023A&A...678A..50E; 2023ApJ...951L...6R; 2023RAA....23g5024X). Direct evidence for such binaries includes 4C+37.11 (2004ApJ...602..123M; 2014ApJ...780..149R; 2017ApJ...843...14B; 2024ApJ...960..110S), whose black holes have projected separation $7.3$ pc, as well as the quasi-periodic blazar OJ 287 (1988ApJ...325..628S; 2022MNRAS.513.3165K). However, despite the promise of searching for periodicity in active galactic nuclei (AGN) light curves as a discovery method, statistically strong SMBHB candidates have proven elusive (2016MNRAS.461.3145V; 2019ApJ...884...36L).
@@
PKS 2131, a BL Lac object at $z_{S}=1.283$, has shown coherent $P=1739.8\pm 17.4$ day periodicity over 47.9 years in radio (2021MNRAS.506.3791R; 2022ApJ...926L..35O; 2024Paper2). PKS J0805, a flat-spectrum radio quasar (FSRQ) at redshift $z_{S}=1.388$ (2007ApJS..171...61H; 2008ApJS..175...97H), has shown a $P=1240.7\pm 4.6$ day periodicity over 14 years of radio observation (2021RAA....21...75R; 2025ApJ...987..191D). Recent observations have revealed the same periodicities across the spectrum, from cm to optical wavelengths, and possibly even gamma-rays (2024Paper2; 2025arXiv250404278H).
Motivated by the observations of PKS 2131 and PKS J0805, we present a model for jetted SMBHB systems. We invoke a sub-relativistic wind to collimate the jet launched from one of the two black holes in the binary. The wind, dragged around by the orbital motion, creates a conical-helical channel through which the relativistic flow travels, producing sinusoidal aberration. [^1] Note that, at radii much greater than the orbital radius, it is the orbital velocity that determines the wind direction. The jet direction consequently varies in quadrature with the displacement and so the jet kinematics is similar to that induced by jet precession. This work enhances the simple model presented in 2022ApJ...926L..35O and 2024Paper2, in which the relativistic jet propagates ballistically, requiring coherent variation over $\sim 100$ pc scales. We detail the model in Sec. 2. We interpret the observations of PKS 2131 and PKS J0805 and apply our model in Sec. 3. In Sec. 4, we consider the prospects of future emission line and GW observations, before concluding in Sec. 5.
@@ 2 Flux variation from Disk-wind Collimated Jets
Figure 1: A visualization of our model. (Top) An SMBHB composed of the primary black hole H and a secondary black hole S. H orbits with speed $v_{H}$, carries around a disk (D) and launches a jet (J) along the orbital angular momentum direction (L). The disk produces a wind (W) with speed $v_{w}<c$, that collimates the jet. (Middle) A geometric illustration of the conical jet along the $\hat{z}$ axis with opening angle $\theta_{j}$ confined by the wind. The wind-jet interface has velocity $\vec{\beta}_{w}=\vec{v}_{w}/c$. We show a unit vector $\hat{r}$ associated with angular position ($\chi$, $\psi$) as well as the observer viewing direction $\hat{n}$. The orbital motion with velocity $\vec{\beta}_{H}=\vec{v}_{H}/c$ will distort the jet into a conical helix. (Bottom) A 2D slice through the wind-collimated helical jet. The dashed line in the center denotes the sinuous path of relativistic flow moving through the center of the channel. The minimum radii $r_{\min}$ which contribute at two frequencies are labeled.
In our model, the SMBHB has period $P$ and is composed of primary mass $M_{1}$ and secondary mass $M_{2}$. We envision one black hole (either the primary or secondary) launching a jet. If the other black hole also launches a jet, we can add a second jet whose orbital motion is half a period out of phase. The main observational feature will be additional higher harmonics. We imagine a sub-relativistic collimating wind, which confines the synchrotron emitting relativistic jet with velocity $\beta_{j}$ and associated bulk Lorentz factor $\Gamma_{j}$. We suppose the interface between the jet and wind has effective speed $\beta_{w}\lesssim 0.9$ (i.e. lying between the sub-relativistic wind and highly relativistic jet speeds). In this model, the wind plays a different role from the sheath in 2017MNRAS.465..161S because it does not contribute to the radiation. The wind serves primarily to confine the jet, while the radiation comes from inside the jet.
The sub-relativistic wind may be a hydromagnetic flow produced from an extended disk (e.g. 1982MNRAS.199..883B; 1992ApJ...385..460E; 2016MNRAS.461.2605G) or a jet sheath (e.g. 1989A&A...224...24P; 1989MNRAS.237..411S; 2017MNRAS.465..161S). We assume the inertia of the wind-jet complex is predominantly carried by the wind. Consequently, we imagine that the relativistic jet direction is completely determined by the wind-jet interface as in the middle panel of Fig. 1. The orbital motion will add to the effective wind velocity, distorting the direction sinusoidally, forming a conical helix as in the bottom panel of Fig. 1. Functionally, this model is similar to those presented in 2017MNRAS.465..161S and 2022ApJ...926L..35O, except that the helix shape and propagation speed are set by the wind-jet interface rather than a ballistically propagating jet.
@@
The inclusion of the wind in our model adds a lever on the emission region locations at different frequencies. If the jet were to propagate ballistically with $\Gamma_{j}$, phase offsets in the observed sinusoidal light curves at different observing frequencies seen in PKS 2131 and PKS J0805 would suggest emission regions separated by $\sim 100$ pc (2024Paper2). These large separations would require the helical structure to remain coherent over $\gtrsim 100$ helical wavelengths $\lambda_{h}=\beta_{w}cP$ (where $\beta_{w}=\beta_{j}$ if the relativistic flow is ballistic). With the jet channel propagating at a non-ultra-relativistic $\beta_{w}$, the emission regions can be separated by $\lesssim 10\lambda_{h}$.
@@ 2.1 Basic Jet Model
We suppose that the jet is confined within a cone of constant opening angle $\theta_{j}$. The observer direction makes an angle $i$ with the cone axis. The jet velocity is radial and constant throughout the jet, so that $\vec{\beta}_{j}=\beta_{j}\hat{r}$, where $\hat{r}$ is a radial unit vector associated with a particular angular position within the open jet (see middle panel of Fig. 1). We assume that the isotropic relativistic electron distribution function in the comoving jet (primed) frame is $n^{\prime}(\gamma)=K^{\prime}\gamma^{-p}$ for $\gamma_{e,\rm min}<\gamma<\gamma_{e,\rm max}$. The associated electron pressure is $P^{\prime}_{e}=\frac{1}{3}K^{\prime}m_{e}c^{2}\int_{\gamma_{e,\mathrm{m}in}}^{\gamma_{e,\mathrm{m}ax}}\gamma^{1-p}d\gamma$, where $m_{e}$ is electron mass. We normalize the pressure using the power in relativistic electrons $L_{e}=4\pi P^{\prime}_{e}(\Gamma_{j}\theta_{j}r)^{2}c$, where $r$ is radius from the black hole. Likewise, the comoving magnetic field strength $B^{\prime}$, assumed to be mostly transverse and disordered, is normalized by the electromagnetic power $L_{m}=(B^{\prime}\Gamma_{j}\theta_{j}r)^{2}c/4$.
@@
The synchrotron emissivity in the lab frame at lab frame frequency $\nu$ and angular position $\Omega$ will be
@@
$\begin{split}j_{\nu,\Omega}=&D^{\frac{3+p}{2}}j^{\prime}_{\nu,\Omega}\\ =&D^{\frac{3+p}{2}}\frac{\sqrt{3\pi}}{16\pi^{2}m_{e}c^{2}}q_{e}^{3}\frac{K^{\prime}}{p+1}\left(\frac{2\pi m_{e}c}{3q_{e}}\right)^{-\frac{p-1}{2}}B^{\prime\frac{p+1}{2}}\\ &\times{\mathrm{\Gamma}}\left(\frac{p}{4}+\frac{19}{12}\right){{\Gamma}}\left(\frac{p}{4}-\frac{1}{12}\right)\\ &\times{{\Gamma}}\left(\frac{p}{4}+\frac{5}{4}\right){\Gamma}\left(\frac{p}{4}+\frac{7}{4}\right)^{-1}\nu^{-\alpha},\end{split}$ (2)
@@
where $j^{\prime}_{\nu,\Omega}$ is the pitch angle-averaged emissivity in the comoving jet frame, $\Gamma$ denotes the Gamma function (not to be confused with jet bulk Lorentz factor $\Gamma_{j}$), $q_{e}$ here is electron charge, and $\alpha=(p-1)/2$ is the spectral index (1979rpa..book.....R). The absorption coefficient in the observer frame is
@@
$\begin{split}\mu_{\nu}=&D^{\frac{p+2}{2}}\mu^{\prime}_{\nu}\\ =&D^{\frac{p+2}{2}}\frac{\sqrt{3\pi}}{32\pi^{2}m_{e}^{2}c^{2}}\left(\frac{3q_{e}}{2\pi m_{e}c}\right)^{\frac{p}{2}}K^{\prime}B^{\prime\frac{p+2}{2}}\\ &\times\Gamma\left(\frac{p}{4}+\frac{22}{12}\right)\Gamma\left(\frac{p}{4}+\frac{2}{12}\right)\\ &\times\Gamma\left(\frac{p}{4}+\frac{6}{4}\right)\Gamma\left(\frac{p}{4}+\frac{8}{4}\right)^{-1}\nu^{-\frac{p+4}{2}},\end{split}$ (3)
@@
where $\mu_{\nu}^{\prime}$ denotes the absorption coefficient in the flow frame. For radio emitting particles in $B\sim 10$ mG fields, the synchrotron cooling times are much longer than the orbital period, so adiabatic losses dominate in the cm emission region. This may not be the case for mm emission regions where the fields may reach up to 1 G, but we leave the inclusion of cooling to future work. For this work, we assume one sustained particle spectrum across the radio emission regions without cooling, which corresponds to assuming a volume-filling acceleration zone with very fast acceleration relative to cooling. We return to acceleration mechanisms in Sec. 3.1.1.
@@
Orbital motion causes major variation in $D$ and consequently flux density variation due to coherent aberration across the jet. To allow coherent helical motion, hydromagnetic waves must be able to cross the jet on a timescale comparable to the flow time through the jet helix (i.e. $\Gamma_{j}\theta_{j}\lesssim 1/\beta_{w}$). In practice, the jet kinematics is likely significantly more complicated and variable. We discuss some of these complications in Sec. 5. The underlying “clock” should not exhibit frequency noise.
@@ 2.2 Kinematics
Due to the orbital motion and collimation by the wind, a radial unit vector $\hat{r}$ in a particular direction within the open jet will be redirected
@@
where $\hat{r}_{0}$ is the unperturbed radial unit vector, which corresponds to a fixed direction in the black hole frame. Note that $\vec{\beta}_{H}$ will introduce a sinusoidal perturbation to $\hat{r}$. At a particular radius $r$ from the black hole, $\vec{\beta}_{H}$ in eq. 4 is delayed in phase from the instantaneous orbital velocity by $r/\lambda_{h}$.
@@ 2.2.1 Amplitude of Flux Density Variations
in the lab frame. We express the line of sight direction as
@@
where $i$ is the angle between $\hat{n}$ and the jet axis at its base (assumed to be along $\hat{z}$). The middle panel of Fig. 1 shows these vectors labeled on the jet.
@@ 2.2.2 Linear Polarization
Given that the magnetic field is assumed to be transverse [^2] Observational justification for our assumption of a toroidal magnetic field is given in 2024Paper2, as the polarization in PKS 2131 is observed to be along the large-scale jet direction., the electric vector should be along $\hat{r}$. The electric vector polarization angle (EVPA) and polarization degree (PD) will vary because the angle between $\hat{r}$ and $\hat{n}$ changes. The polarization properties discussed here will be for one specific position ($r,\chi,\psi$) in the jet. Different $(r,\chi,\psi)$ positions will have the same general polarization pattern but be slightly out of phase, leading to some decoherence.
@@
$\vec{e}_{obs}=\frac{\hat{r}-\left(\hat{r}\cdot\hat{n}\right)\hat{n}}{\sqrt{1-\left(\hat{r}\cdot\hat{n}\right)^{2}}}.$ (18)
The observed PD will be a function of viewing angle. For a transverse field, more edge on vantage points give higher polarization. Thus, the observed PD is
$\Pi_{\rm obs}=\Pi_{\max}\left[1-\cos^{2}\theta\right]=\Pi_{\max}\left[1-\left(\hat{r}\cdot\hat{n}\right)^{2}\right],$ (19)
where $\Pi_{\max}$ is the PD when viewing completely edge on. $\Pi_{\max}$ will be suppressed from the nominal synchrotron polarization limit due to the level of magnetic turbulence present in the emission zone.
We decompose eq. 18 and eq. 19 into a constant and varying component